# Manturov Projection for Virtual Legendrian Knots in $ST^{*}F$

### v1

@@ Abstract.
We show that there is a projection operation $\mathop{\mathrm{proj}}\nolimits$ from the set of virtual isotopy classes of Legendrian knots to the set of isotopy classes of Legendrian knots in $ST^{*}S^{2}$. This projection is obtained by substituting some of the classical crossings of the front diagram for a virtual crossing. It restricts to the identity map on the set of virtual isotopy classes of classical Legendrian knots. In particular, the projection $\mathop{\mathrm{proj}}\nolimits$ extends invariants of Legendrian knots to invariants of virtual Legendrian knots. Using the projection $\mathop{\mathrm{proj}}\nolimits$, we show that the virtual crossing number of every classical Legendrian knot equals its crossing number. We also prove that the virtual canonical genus of a Legendrian knot is equal to the canonical genus.
@@ 1. Introduction
As proved by Goussarov, Polyak, and Viro [13] (see also Kauffman [14]), two different classes of knots do not become equivalent in the virtual category. Carter, Kamada, Saito [5] proved that virtual knots can be viewed as equivalence classes of knots in a thickened surface $F\times\mathbb{R}$ modulo isotopies and stabilizations/destabilizations of $F$ (adding and deleting handles) away from the knot projection, as well as automorphisms of $F\times\mathbb{R}$ of the form $\alpha\times\mathop{\mathrm{id}}\nolimits_{\mathbb{R}}$ where $\alpha$ is an orientation preserving diffeomorphism of $F$.
@@
A Legendrian submanifold $L\subset ST^{*}M$ is fully described by its naturally cooriented front projection $\pi(L)\subset M$. Note that in general the front of $L$ is not immersed and has various Legendrian singularities such as swallow tails etc.
@@
2. Projection and Parities for Virtual Knots in $ST^{*}F$
This section is largely motivated by the ideas of the Manturov projection [21] of virtual knots in $F\times\mathbb{R}.$
Consider an irreducible Legendrian knot $L$ in $ST^{*}F$. Let $D(L)$ be the diagram of $L$. We will denote the classical intersection points of $D(L)$ by the same letters as the intersection points of the front diagram $\pi(L)$ of $L$. Each classical self-crossing $p$ of $\pi(L)$ separates $\pi(L)$ into two loops that are obtained from $\pi(L)$ by choosing an orientation of $L$, and then smoothing $\pi(L)$ at $p$ according to the orientation of $L$. We call the self-crossing $p$ of $\pi(L)$ odd if each crossing $p$ of $\pi(L)$ separates it into two loops and if it is not true that the classes of both loops in $\pi_{1}(F,p)$ are powers of the same third element, then $p$ is odd. Note that these powers can be positive, zero or negative.
Manturov projection of a virtual irreducible Legendrian knot $L$ in $ST^{*}F$ is the classical Legendrian knot $\mathop{\mathrm{proj}}\nolimits(L)$ such that the diagram $D(\mathop{\mathrm{proj}}\nolimits(L))$ is obtained from $D(L)$ by changing each odd crossing of $\pi(L)$ to a virtual crossing and keeping even crossings of $\pi(L)$ intact. If $L$ is a virtual Legendrian knot that is not necessarily irreducible, then, by our generalization [11, Theorem 1.1] of the Kuperberg [16] Theorem, virtual isotopy class of $L$ has a unique irreducible representative $L^{\prime}$, defined up to Legendrian isotopy and automorphism of $F$. We define $\mathop{\mathrm{proj}}\nolimits(L)$ to be $\mathop{\mathrm{proj}}\nolimits(L^{\prime})$.
Theorem 2.1.
The Legendrian isotopy class of the Manturov projection of a virtual Legenrian knot in $ST^{*}F$ does not depend on the representative of a virtual knot and maps virtually Legendrian isotopic knots to isotopic Legendrian knots. In particular, if a virtual Legendrian knot $L_{v}$ is virtual isotopic to a classical Legendrian knot $L$ in $ST^{*}S^{2}$, then its projection $\mathop{\mathrm{proj}}\nolimits(L_{v})$ is Legendrian isotopic to that knot $L$.
The proof of Theorem 2.1 follows immediately from Theorem 3.1.
3. Nikonov theorem
Theorem 3.1.
There is a map $pr$ from diagrams of irreducible virtual Legendrian knots to classical knot diagrams such that for every knot $K$ we have $\mathop{\mathrm{proj}}\nolimits(K)$ is obtained from $K$ by substituting odd crossings to virtual crossings; and if two diagrams $K_{1}$ and $K_{2}$ are related by a Reidemeister move of Legendrian knots (performed within the given minimal genus diagram), then their images $\mathop{\mathrm{proj}}\nolimits(K_{1})$ and $\mathop{\mathrm{proj}}\nolimits(K_{2})$ are related by a Reidemeister move of Legendrian knots.
Proof.
Let $K$ be an irreducible Legendrian knot in $ST^{*}F$ for a surface $F$ of genus $g$. Recall that we denote by $\pi$ the projection of the circle bundle $ST^{*}F$ onto the base surface $F$. If the front $\pi(K)$ of $K$ is generic, then it only has transverse double points and semi-cubical cusps. By identifying $ST^{*}F$ with the set of cooriented contact elements at points of $F$, we deduce a coorientation of the front $\pi(K)$. Let $D(K)$ denote the cooriented diagram in $\mathbb{R}^{2}$ of the virtual Legendrian knot $K$. The diagram $D(K)$ can be obtained from $K$ by the composition of the projection of $K$ to $F\setminus\{pt\}$ for some point $pt$ outside of $\pi(K)$, and an immersion $F\setminus\{pt\}\to\mathbb{R}^{2}$. All crossings of $D(K)$ that do not correspond to crossings of $\pi(K)$ are marked as virtual. We refer to non-virtual crossings as to classical.
Every classical crossing $v$ corresponds to two loops $h_{v,1}$ and $h_{v_{2}}$ of $\pi(K)$ on $F$. We define $\mathop{\mathrm{proj}}\nolimits$ by replacing some of the classical crossings of $D(K)$ with virtual ones. Namely, if the homotopy classes of $h_{v,1}$ and $h_{v,2}$ are powers of the same element of the fundamental group, then the crossing remain classical, and otherwise it becomes virtual. Here we permit the power to be positive, negative, or zero. As in the case of virtual knots, $\mathop{\mathrm{proj}}\nolimits(K)$ is a diagram of a Legendrian knot over the sphere $S^{2}$. Thus, indeed there is a map $\mathop{\mathrm{proj}}\nolimits$.
Suppose that the knot $K^{\prime}$ is obtained from the knot $K$ by a cusp birth move, which corresponds to the first Reidemeister move of knots. Then at least one of the two loops, $h_{v,1}$ or $h_{v,2}$, is trivial. Therefore, the crossing remains classical, which means that $\mathop{\mathrm{proj}}\nolimits(K^{\prime})$ is obtained from $\mathop{\mathrm{proj}}\nolimits(K)$ by the same move.
The second Reidemeister move of knots corresponds to the fold-cusp move (when a cusp pinches a branch) and to the two possible safe self-tangencies when the coorientations of the tangent branches point in opposite directions and hence do not give a double point of $K.$ As in the argument for knots, either the two new crossings in $K^{\prime}$ remain in $\mathop{\mathrm{proj}}\nolimits(K^{\prime})$, or both new crossings in $K^{\prime}$ become virtual in $\mathop{\mathrm{proj}}\nolimits(K^{\prime})$.
The third Reidemeister move of Legendrian knots is similar to the third Reidemeister move for knots. By an argument in the proof of [21, Theorem 6], out of three crossings in $K$ either only one remains classical, or all three remain classical in $\mathop{\mathrm{proj}}\nolimits(K)$. (So the case where exactly one of the crossings becomes virtual is not possible.) If all three crossings remain classical, then $\mathop{\mathrm{proj}}\nolimits(K^{\prime})$ is obtained from $\mathop{\mathrm{proj}}\nolimits(K)$ by the same third Reidemeister move.
Suppose now that only one of the three crossings in $K$ remain classical in $\mathop{\mathrm{proj}}\nolimits(K)$. Then in $\mathop{\mathrm{proj}}\nolimits(K^{\prime})$ also only one of the crossings in classical. Thus, again $\mathop{\mathrm{proj}}\nolimits(K^{\prime})$ is obtained from $\mathop{\mathrm{proj}}\nolimits(K)$ by the third Reidemeister move for virtual knots. ∎
@@ 4. Crossing number and canonical genus corollaries of the Projection Theorem
Given a knot diagram $D$ the knot $K$ its canonical genus $g(D)$ is the genus of a surface obtained as follows: smoothen the classical crossings of $D$ using orientation of the knot branches and glue the resulting ovals with disks, then glue the bands to the disks at each crossing. The canonical genus $\mathop{\mathrm{cg}}\nolimits(K)$ is the minimum of $g(D)$ over all the knot diagrams of $D.$ Clearly this procedure gives also the virtual canonical genus $\mathop{\mathrm{vcg}}\nolimits(K_{v})$ of a virtual knot $K_{v}$ with the minimum taken over the class of virtual knot diagrams of $K_{v}.$ It is clear that when the virtual knot $K_{v}$ is isotopic to a classical knot $K$ we have $\mathop{\mathrm{vcg}}\nolimits(K_{v})\leq\mathop{\mathrm{cg}}\nolimits(K)$, see Stoimenov, Tchernov, Vdovina [26]. In [6] the first author used the Manturov Projection Theorem to prove that for a classical knot $K$ we have $\mathop{\mathrm{vcg}}\nolimits(K)=\mathop{\mathrm{cg}}\nolimits(K)$.

### v2

@@ Abstract.
We show that there is a projection operation $\mathop{\mathrm{proj}}\nolimits$ from the set of virtual isotopy classes of Legendrian knots to the set of isotopy classes of Legendrian knots in $ST^{*}S^{2}$. This projection is obtained by substituting some of the classical crossings of the front diagram with virtual crossings. It restricts to the identity map on the set of virtual isotopy classes of classical Legendrian knots. In particular, the projection $\mathop{\mathrm{proj}}\nolimits$ extends invariants of Legendrian knots to invariants of virtual Legendrian knots. Using the projection $\mathop{\mathrm{proj}}\nolimits$, we show that the virtual crossing number of every classical Legendrian knot equals its crossing number. We also prove that the virtual canonical genus of a Legendrian knot is equal to the canonical genus.
@@ 1. Introduction
As proved by Goussarov, Polyak, and Viro [13] (see also Kauffman [15]), two non isotopic knots do not become equivalent in the virtual category. Carter, Kamada, Saito [5] proved that virtual knots can be viewed as equivalence classes of knots in a thickened surface $F\times\mathbb{R}$ modulo isotopies and stabilizations/destabilizations of $F$ (adding and deleting handles) away from the knot projection, as well as automorphisms of $F\times\mathbb{R}$ of the form $\alpha\times\mathop{\mathrm{id}}\nolimits_{\mathbb{R}}$ where $\alpha$ is an orientation preserving diffeomorphism of $F$.
@@
A Legendrian submanifold $L\subset ST^{*}M$ is fully described by its naturally cooriented front projection $\pi(L)\subset M$. Note that in general the front of $L$ is not immersed and has various Legendrian singularities such as cusps.
@@ 2. Crossing number and canonical genus corollaries of the Projection Theorem
In section 4 we will prove the Projection Theorem (Theorem 4.1) for virtual Legendrian knots asserting that there is a map $upr$ that associates a classical Legendrian knot diagram in $\mathbb{R}^{2}$ to each virtual Legendrian knot diagram in $\mathbb{R}^{2}$. Furthermore, the generalized Gauss diagram of the Legendrian knot $upr(K)$ is obtained from the generalized Gauss diagram of the virtual Legendrian knot $K$ by erasing some of the chords and leaving the other chords intact, and if $K$ is virtual Legendrian isotopic to a classical Legendrian knot, then $upr(K)$ belongs to the virtual Legendrian isotopy class of $K$.
In this section we will use the Projection Theorem to deduce two consequences, namely, that the virtual crossing number of every classical Legendrian knot equals its crossing number, and that the virtual canonical genus of a Legendrian knot coincides with the canonical genus.
@@
Given a knot diagram $D$ of the knot $K$ its canonical genus $g(D)$ is the genus of a canonical surface $S$ obtained as follows: smoothen the classical crossings of $D$ using orientation of the knot branches and glue the resulting ovals with disks. Then glue the bands to the disks at each crossing. There is an obvious copy $K^{\prime}$ of the knot $K$ over the surface $S$. We say that $(K^{\prime},S)$ is a canonical representative of the knot diagram $D$. The canonical genus $\mathop{\mathrm{cg}}\nolimits(K)$ is the minimum of $g(D)$ over all the knot diagrams of $D.$ Clearly this procedure gives also the virtual canonical genus $\mathop{\mathrm{vcg}}\nolimits(K_{v})$ of a virtual knot $K_{v}$ with the minimum taken over the class of virtual knot diagrams of $K_{v}.$ It is clear that when the virtual knot $K_{v}$ is isotopic to a classical knot $K$ we have $\mathop{\mathrm{vcg}}\nolimits(K_{v})\leq\mathop{\mathrm{cg}}\nolimits(K)$, see Stoimenov, Tchernov, Vdovina [27]. In [6] the first author used the Manturov Projection Theorem to prove that for a classical knot $K$ we have $\mathop{\mathrm{vcg}}\nolimits(K)=\mathop{\mathrm{cg}}\nolimits(K)$.